OpenStudy (anonymous):

Seven times a number is 9 less than the product of the number and -20. What is the number

OpenStudy (texaschic101):

7n = -20n - 9 --- add 20n to both sides 7n + 20n = -9 27n = -9 --- divide by 27 n = -9/27 reduce n = - 1/3

OpenStudy (anonymous):

Thank you @texaschic101